find three consecutive numbers whose sum is 15 more than twice the smallest number
Answer by mananth(16949)

find three consecutive numbers whose sum is 15 more than twice the smallest number
...
let the numbers be x,x+1,x+2
..
x+x+1+x+2=2x+15
3x+3=2x+15
-2x
3x-2x+3=15
x+3=15
-3
x=12
12,13,14
..
m.ananth@hotmail.ca